10)A computer program is tested by 5 independent tests. If there is an error, these tests will discover it with probabilities 0.1, 0.2, 0.3, 0.4, and 0.5, respectively. Suppose that the program contains an error. What is the probability that it will be found
g. by at least one test?
h. by at least two tests?
i. by all five tests?

Answer by robertb(5830)

g. 1 - 0.9*0.8*0.7*0.6*0.5 = 0.8488
h. 1 - 0.9*0.8*0.7*0.6*0.5 - 0.1*0.8*0.7*0.6*0.5 - 0.9*0.2*0.7*0.6*0.5 - 0.9*0.8*0.3*0.6*0.5 - 0.9*0.8*0.7*0.4*0.5 - 0.9*0.8*0.7*0.6*0.5 = 0.8488 - 0.0168-0.0378-0.0648-0.1008 - 0.1512 = 0.4774
i. 0.1*0.2*0.3*0.4*0.5 = 0.0012